# Relating Checkpoint Update Probabilities to Momentum Parameters in Single-Loop Variance Reduction Methods

### v1

@@ 2 Notations and Assumptions
We define computational costs by making use of the IFO framework of [16], where we assume that sampling an index $i$ and computing the pair $(f_{i}(x),\nabla f_{i}(x))$ incurs a unit of cost. In the composite general convex setting, $x$ is called an $\epsilon$-accurate solution iff $\mathbb{E}\left\{F(x)-F(x^{*})\right\}\leq\epsilon$, where $x^{*}$ is a solution to the problem (1). The IFO complexity is the minimum IFO calls needed to reach an $\epsilon$-accurate solution.
@@
A4 The function $l$ is convex, lower semi-continuous (but possibly non-differentiable) function which admits an efficient proximal operator.
3 The Katyusha-H Method
The proposed Katyusha-H method is formally described in Algorithm 1. The three sequences formulation, consisting of ${x_{t}},{y_{t}},{z_{t}}$, is ubiquitous in accelerated gradient methods [37] and can be interpreted by the linear coupling interpretation [38] as follows. Consider a simplified setting where $l\equiv 0$ so that $F\equiv f$ and the mini-batch size $b=n$ so that $\tilde{g}_{t+1}=\nabla f(x_{t+1})$. Then the $\{z_{t}\}$ sequence is viewed and analysed as a mirror descent sequence, where an error term proportional to $\|\nabla f(x_{t+1})\|^{2}$ is introduced in its analysis. The $\{y_{t}\}$ iteration can be equivalently formulated as $y_{t+1}=y_{t}-\tau_{t}\alpha_{t}\eta\nabla f(x_{t+1})$, which further reduces to $y_{t+1}=y_{t}-\eta\nabla f(x_{t+1})$ when $\tau_{t}\alpha_{t}=1$. Hence $\{y_{t}\}$ is viewed and analysed as a gradient descent sequence, where an improvement proportional to $\|\nabla f(x_{t+1})\|^{2}$ is made in each iteration. Finally, $x_{t+1}$ linearly couples $z_{t}$ and $y_{t}$ to achieve acceleration, in the spirit of leveraging the aforementioned duality of the mirror descent and the gradient descent.
@@ 3 The Katyusha-H Method
Algorithm 1 Katyusha-H
@@
The checkpoint $w_{t}$ is updated in a probabilistic way, which was first proposed by [9]. Most single-loop VR methods, including the pioneering work [9], use constant checkpoint update probability. To the best of our knowledge, the SVRLS method [39] is the only method that uses varying update probability $p_{t}=1/(at+1)$ where $0<a<1$. However, SVRLS is a non-accelerated method that is analysed by an AdaGrad-type regret analysis under an extra compact feasible set condition, requires that the nonsmooth part $l\equiv 0$ in (1), and unfortunately fails to provide convergence guarantee for the iterate sequence itself [^5] SVRLS thus does not have the last-iterate convergence guarantee..
The key feature of Algorithm 1 is the utilization of the proposed Harmonia technique, which is instrumental to the theoretical improvement of this work, and also leads to some novel insights into single-loop VR methods. These will be discussed concretely in the next section, where we present theoretical properties of the Katyusha-H method.
@@ 4 Main Results and Related Discussions
and $\tilde{\alpha}_{0}=\xi\alpha_{1}^{2}$. Then in Algorithm 1, $p_{t}\in[0,1]$ for $t\in\mathbb{N}^{+}$, $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq 0$ for $t\in\mathbb{N}$, and $\tau_{t},\xi,(1-\xi-\tau_{t})\in(0,1)$ for $t\in\mathbb{N}^{+}$. Moreover, there exists a constant $C>0$ which does not depend on any problem-dependent constants (e.g., $n$) such that [^6] As is shown in the proof, $C$ can take a mild value, e.g., $C=5$.
@@
The setting of parameters of Algorithm 1 are determined in Lemma 1 and Theorem 1, and it is fairly intuitive. First, $\xi=1/(bc)$ is the coefficient of the Katyusha momentum, which is constructed to cancel certain terms in the upper bound of the variance of the SVRG estimator, i.e., to further reduce the variance (in the sense that one can use a tighter upper bound on the variance of the SVRG estimator with the presence of the Katyusha momentum). Hence it is expected that $\xi$ becomes smaller when the variance of the SVRG estimator decreases, which is actually the case since $\xi$ is non-increasing (and likely to decrease) when the mini-batch size $b$ increases. Second, the largest allowable step size $\eta=1/(cL+L)$ is non-decreasing (and likely to increase) when $b$ increases, which is reasonable since larger step size is expected to work when the variance of the stochastic estimator decreases.
Note that the convergence guarantee is for the iterate sequence $\{w_{t}\}$ itself [^7] By which we mean $F(w_{t})$ converges to $F(x^{*})$ in expectation. Convergence guarantee is also available for the $\{y_{t}\}$ sequence. but the convergence rate of $\{y_{t}\}$ is slower than that of $\{w_{t}\}$., hence the last-iterate convergence is available. In the right hand side (RHS) of (2), the coefficient $\alpha_{0}^{2}$ equals $36$ and is an absolute constant, $\tilde{\alpha}_{0}$ is upper bounded by $36$, and $1/(2\eta)$ equals $L(c+1)/2$ when taking the largest allowable step size, where $c\leq C$ by Lemma 1. Hence the convergence rate is $\mathcal{O}(1/(\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}))$ and is $n$-independent [^8] We refer to the proof of Theorem 2 in the appendix, where we show that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}$ is lower bounded by a constant multiple of $T^{\alpha+1}$., which is in contrast to Theorem $5$ in [30], where the convergence rate has an $n$-dependent constant and the order of $n$ can be very large [^9] In Theorem $5$ of [30], using the same notation as [30], the dominant constant of the convergence rate is either $\mathcal{O}(\nu^{2})$ or $\mathcal{O}(c)$, which are both $n$-dependent..
It is interesting to notice the different roles of the checkpoint update probability $p_{t}$ and the mini-batch size $b$ in the convergence analysis. In expectation, one iteration of Algorithm $1$ costs $\mathcal{O}(np_{t}+b)$ IFO calls, hence $p_{t}$ and $b$ both relate to the per-iteration cost. In theory, $p_{t}$ and $b$ both relate to the variance of the SVRG estimator. When $p_{t}\equiv 1$ or $b=n$, the variance vanishes and the SVRG estimator reduces to the full gradient. However, in the convergence analysis, $p_{t}$ is chosen in accordance with $\alpha_{t}$ while $b$ can be chosen arbitrarily. This observation indicates that $p_{t}$ plays a more significant role in maintaining the convergence property.
@@
Under the setting of Lemma 1 and Theorem 1, for any fixed $\alpha\in[0,1]$ in the definition of $\{\alpha_{t}\}_{t\in\mathbb{N}}$ in Lemma 1, the convergence rate of Algorithm 1 is $\mathcal{O}(1/T^{\alpha+1})$, the per-iteration IFO call in expectation is $\mathcal{O}(b+n/(bt^{1-\alpha})+n/t)$ for all $t\geq 17$. Without loss of generality, assume that the target accuracy $\epsilon\in(0,1)$, define $\hat{\alpha}=\log(2)/\log(\lceil 1/\epsilon\rceil)$, then the IFO call in expectation needed by Algorithm 1 to reach an $\epsilon$-accurate solution is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left[0,\min\left\{\hat{\alpha},1/10\right\}\right]\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\frac{n}{b}\frac{1}{\alpha}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\log(1/\epsilon)\right).\quad&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right]\end{cases}$ (3)
@@
We give some remarks on the threshold $\hat{\alpha}$. By the proof of Theorem 2, the complexity of the $\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right]$ case actually applies to all $\alpha\in\left(0,1\right]$. But to prevent the constant $1/\alpha$ from exploding when $\alpha\rightarrow 0^{+}$, more delicate proof is conducted where we introduce the threshold $\hat{\alpha}$ and discuss the two cases separately and carefully.
@@
In the finite-sum optimization, one of the most fundamental topics is the trade-off between the iteration complexity and the per-iteration IFO cost, since essentially it is these two factors that determines the theoretical complexity of an optimization method. Handling this trade-off typically involves using different gradient estimators, using different optimization techniques, and even changing the algorithmic structure. For example, using full gradient gives the proximal gradient descent (PGD) a fast iteration convergence rate and high per-iteration IFO cost. While using the SGD estimator gives the proximal stochastic gradient descent (PSGD) contradictory characteristics. On the algorithmic structure side, note that the majority of variance reduction methods are nested, while PGD and PSGD are single-loop methods.
However, based on the proposed $p_{t}-\alpha_{t}$ correspondence, we can achieve this trade-off under the unifying framework of Algorithm 1, by changing the growth rates of the $\{\alpha_{t}\}_{t\in\mathbb{N}}$ sequence, which is novel to the literature to the best of our knowledge.
@@
As proved in Theorem 2, less aggressive momentum acceleration (i.e., smaller $\alpha$) accompanies with smaller checkpoint update probability (i.e., smaller $p_{t}$), hence slower convergence rate can be compensated with lower per-iteration IFO cost in Algorithm 1, showing a novel continuous [^10] By continuous, we mean that the parameter $\alpha$ is allowed to take any value in the interval $[0,1]$. trade-off between the convergence rate and the per-iteration cost. The key parameter $\alpha\in[0,1]$ is thus a weight that controls the proposed acceleration-variance reduction trade-off, through which Algorithm 1 essentially controls the trade-off between the convergence rate and the per-iteration cost. Two consequences of this mechanism are listed as follows.
@@
First, it is interesting to note that Algorithm 1 recovers some existing IFO complexity of very different algorithms under a unifying framework. Consider the $b=1$ case. When $\alpha=0$ (i.e., $\alpha_{t}\equiv 6$ for all $t\in\mathbb{N}$), the complexity of Algorithm 1 is $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$, recovering the complexity of SVRG++ while removing some requirements concerning the knowledge of the solution $x^{*}$ needed by SVRG++. When $\alpha=1$ (i.e., $\alpha_{t}=\mathcal{O}(t)$), the complexity of Algorithm 1 is $\mathcal{O}(n+n/\sqrt{\epsilon})$, recovering the complexity of FISTA. Note that SVRG++ is a nested algorithm, uses the SVRG estimator, and employs the doubling technique, while FISTA is a single-loop algorithm, uses the full gradient as the iteration direction, and employs the Nesterov acceleration technique.
@@
Second, since $\alpha$ in Lemma 1 belongs to $[0,1]$, we discover a series of novel complexity, from which one can choose a suitable $\alpha$ value to obtain near-optimal method according to the knowledge of $n$ and $\epsilon$ under a mild condition. Consider the $b=1$ case for Algorithm 1. When $n=10^{4},\epsilon=10^{-12}$, the complexity of SVRG++, FISTA, Algorithm 1 ($\alpha=1/2$), and the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ are $\mathcal{O}(10^{12})$, $\mathcal{O}(10^{10})$, $\mathcal{O}(10^{8})$ and $\mathcal{O}(10^{8})$, respectively. In fact, given $n\in\mathbb{N}^{+}$ and $\epsilon\in(0,1)$ satisfying
@@
Hence the complexity in expectation of Algorithm 1 matches the theoretical lower bound up to a logarithmic factor under the $\alpha$ setting (5). This observation can be summarised in the following proposition.
@@
Under the setting and assumptions of Theorem 2, assume that condition (4) holds, by choosing $b=1$ and $\alpha$ satisfying (5), the complexity in expectation of Algorithm 1 is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
@@
Since $\Delta_{1}<\Delta_{2}$ and $\Delta_{2}>0$ under the condition (4) and the appropriate $C_{2}$ choice described in the footnote, the interval $\Delta$ is well-defined, and the intervals in (5) are well-defined and contained in $[0,1]$. Direct calculation shows that $\alpha\geq\Delta_{1}$ implies $1/\epsilon^{1/(\alpha+1)}\leq C_{1}\sqrt{n}/\sqrt{\epsilon}$ and $\alpha\leq\Delta_{2}$ implies $n/\epsilon^{\alpha/(\alpha+1)}\leq C_{2}\sqrt{n}/\sqrt{\epsilon}$. If $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$, under the first choice in (5), we have $\mathbb{E}C_{\text{comp}}$ takes the first form in (3) and $1/\epsilon^{1/(\alpha+1)}\leq\sqrt{n}/\sqrt{\epsilon}$, hence $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. Similarly, one sees that if $\Delta_{2}>\min\left\{\hat{\alpha},1/10\right\}$, we have $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under the second $\alpha$ choice in (5). In all, under the setting of this corollary, $\alpha$ satisfying (5) always exists and such choice of $\alpha$ yields the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. ∎
@@
For ease of discussion, we take $C_{2}=1$ in Corollary 1 and assume that $n$ in (1) is given. Some comments on Corollary 1 are listed as follows. First, we give intuitive explanations for the choice (5). The largest possible interval in (5) under the $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$ case is $\left[0,\min\left\{\hat{\alpha},1/10\right\}\right]$, and the largest possible interval in (5) under the other case is $\left[\min\left\{\hat{\alpha},1/10\right\},1\right]$, which is separated from the former interval by $\min\left\{\hat{\alpha},1/10\right\}$. Since $n$ is given and $C_{2}=1$, $\Delta_{2}$ is smaller when $\log(n)/\log(1/\epsilon)$ is larger, i.e., $\epsilon$ is larger. Hence the selection rule (5) essentially suggests that we should use mild acceleration and small update probability when the target accuracy is low (i.e., $\epsilon$ is relatively large), and we should use aggressive acceleration and large update probability when the target accuracy is high.
For ease of understanding, note that when the update probability $p_{t}$ is small, the per-iteration IFO cost is low and the variance of the gradient estimator $\tilde{g}_{t+1}$ is high, hence we refer to Algorithm 1 under small $p_{t}$ as the stochastic regime. Similarly, we refer to Algorithm 1 under large $p_{t}$ as the pseudo-deterministic regime [^12] Equivalently, Algorithm 1 with small $\alpha$ is the stochastic regime, and it is the pseudo-deterministic regime when $\alpha$ is large.. To understand the proper choice of regimes under different scales of the target accuracy, we turn to the following analogy.
Recalling the familiar algorithms PSGD and PGD, it is well known that PSGD is more suitable for solving (1) when the target accuracy is low (i.e., $\epsilon$ is large), since PSGD has low per-iteration cost and can make rapid progress in the early stage of optimization. However, when $\epsilon$ is small, PSGD has bad $\epsilon$ dependence in the complexity and converges very slowly, which is essentially due to the non-diminishing variance of the gradient estimator. Hence PGD can be favourable for solving (1) to a high target accuracy.
By analogy, for Algorithm 1, it is fairly reasonable to use the stochastic regime when $\epsilon$ is large, and use the pseudo-deterministic regime when $\epsilon$ is small. Moreover, it is well-known that the variance of the gradient estimator should be reduced to certain degree to make acceleration technique effective in randomized algorithms [^13] For example, [40, 41] use a fairly large mini-batch to make the Nesterov acceleration effective. For another example, the Katyusha momentum is proposed to cancel certain terms that come from a tight estimate of the variance of the SVRG estimator, so that the Nesterov acceleration technique can be utilized to accelerate SVRG in terms of the IFO complexity.. Hence, in the stochastic regime, mild acceleration can be added since the variance of the gradient estimator is mildly reduced, while aggressive acceleration can be applied in the pseudo-deterministic regime since the variance is significantly reduced.
Quantitatively, the important issue is exactly how much acceleration (controlled by $\alpha_{t}$) can the method adopt, given certain level of variance reduction (controlled by $p_{t}$). The answer to this question is provided by the proposed $p_{t}-\alpha_{t}$ correspondence, which essentially determines how much variance reduction should be conduct to perform certain level of acceleration, or vise versa.
For a clear interpretation, in the above analysis, the acceleration side and the variance reduction side is separated, and acceleration is regarded as an optional choice based on the existing level of variance reduction. While in Algorithm 1, the correspondence of the strength of acceleration and that of variance reduction comes simultaneously through the $p_{t}-\alpha_{t}$ correspondence, hence the name Harmonia [^14] Harmonia is the name of the Greek goddess of harmony and concord. for the proposed technique.

### v2

@@ 2 Notations and Assumptions
We define computational costs by making use of the IFO framework of [agarwal2015lower], where we assume that sampling an index $i$ and computing the pair $(f_{i}(x),\nabla f_{i}(x))$ incurs a unit of cost. In the composite general convex setting, $x$ is called an $\epsilon$-accurate solution iff $\mathbb{E}\left\{F(x)-F(x^{*})\right\}\leq\epsilon$, where $x^{*}$ is a solution to the problem (1). The IFO complexity is the minimum IFO calls needed to reach an $\epsilon$-accurate solution and is denoted by $C_{comp}$.
@@
A4 The function $l$ is convex, lower semi-continuous (but possibly non-differentiable) which admits an efficient proximal operator.
3 The Acceleration Framework
The proposed framework is formally described in Algorithm 1. The three sequences formulation, consisting of ${x_{t}},{y_{t}},{z_{t}}$, is ubiquitous in accelerated gradient methods [d2021acceleration] and can be interpreted by the linear coupling interpretation [allen2017linear] as follows. Consider a simplified setting where $l\equiv 0$ so that $F\equiv f$ and the mini-batch size $b=n$ so that $\tilde{g}_{t+1}=\nabla f(x_{t+1})$. Then the $\{z_{t}\}$ sequence is viewed and analysed as a mirror descent sequence, where an error term proportional to $\|\nabla f(x_{t+1})\|^{2}$ is introduced in its analysis. The $\{y_{t}\}$ iteration can be equivalently formulated as $y_{t+1}=y_{t}-\tau_{t}\alpha_{t}\eta\nabla f(x_{t+1})$, which further reduces to $y_{t+1}=y_{t}-\eta\nabla f(x_{t+1})$ when $\tau_{t}\alpha_{t}=1$. Hence $\{y_{t}\}$ is viewed and analysed as a gradient descent sequence, where an improvement proportional to $\|\nabla f(x_{t+1})\|^{2}$ is made in each iteration. Finally, $x_{t+1}$ linearly couples $z_{t}$ and $y_{t}$ to achieve acceleration, in the spirit of leveraging the aforementioned duality of the mirror descent and the gradient descent.
@@ 3 The Acceleration Framework
Algorithm 1 The single-loop variance-reduced acceleration framework
@@
The checkpoint $w_{t}$ is updated in a probabilistic way, which was first proposed by [kovalev2020don]. Most single-loop VR methods, including the pioneering work [kovalev2020don], use constant checkpoint update probability. To the best of our knowledge, the SVRLS method [jiang2023adaptive] is the only method that uses varying update probability $p_{t}=1/(at+1)$ where $0<a<1$. However, SVRLS is a non-accelerated method that is analysed by an AdaGrad-type regret analysis under an extra compact feasible set condition, requires that the nonsmooth part $l\equiv 0$ in (1), and unfortunately fails to provide convergence guarantee for the iterate sequence itself [^3] SVRLS thus does not have the last-iterate convergence guarantee..
The key feature of Algorithm 1 is the utilization of the proposed $p_{t}-\alpha_{t}$ correspondence, which is instrumental to the theoretical improvement of this work, and also leads to some novel insights into single-loop VR methods. These will be discussed concretely in the next section, where we present theoretical properties of Algorithm 1.
@@ 4 Main Results and Related Discussions
and $\tilde{\alpha}_{0}\geq\xi\alpha_{1}^{2}$. Then in Algorithm 1, $p_{t}\in[0,1]$ for $t\in\mathbb{N}^{+}$, $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq 0$ for $t\in\mathbb{N}$, and $\tau_{t},\xi,(1-\xi-\tau_{t})\in(0,1)$ for $t\in\mathbb{N}^{+}$. Moreover, there exists a constant $C>0$ which does not depend on any problem-dependent constants (e.g., $n$) such that [^4] As is shown in the proof, $C$ can take a mild value, e.g., $C=5$.
@@
The setting of parameters of Algorithm 1 are determined in Lemma 1 and Theorem 1, and it is fairly intuitive. First, $\xi=1/(bc)$ is the coefficient of the Katyusha momentum, which is constructed to cancel certain terms in the upper bound of the variance of the SVRG estimator, i.e., to further reduce the variance (in the sense that one can use a tighter upper bound on the variance of the SVRG estimator with the presence of the Katyusha momentum). Hence it is expected that $\xi$ becomes smaller when the variance of the SVRG estimator decreases, which is actually the case since $\xi$ is non-increasing (and likely to decrease) when the mini-batch size $b$ increases. Second, the largest allowable step size $\eta=1/(cL+L)$ is non-decreasing (and likely to increase) when $b$ increases by the definition of $c$, which is reasonable since larger step size is expected to work when the variance of the stochastic estimator decreases.
Note that the convergence guarantee is for the iterate sequence $\{w_{t}\}$ itself [^5] By which we mean $F(w_{t})$ converges to $F(x^{*})$ in expectation., hence the last-iterate convergence is available. In the right hand side (RHS) of (2), the coefficient $\alpha_{0}^{2}$ equals $36$ and is an absolute constant, $\tilde{\alpha}_{0}$ can also be set to an absolute constant, and $1/(2\eta)$ equals $L(c+1)/2$ when taking the largest allowable step size, where $c\leq C$ by Lemma 1. Hence the convergence rate is $\mathcal{O}(1/(\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}))$ and is $n$-independent [^6] In Appendix D, we show that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{T}^{2}+\sum_{j=1}^{T}\alpha_{j}$ is lower bounded by a constant multiple of $T^{\alpha+1}$., which is in contrast to Theorem $5$ in [driggs2022accelerating], where the convergence rate has an $n$-dependent constant and the order of $n$ can be very large [^7] In Theorem $5$ of [driggs2022accelerating], using the same notation as [driggs2022accelerating], the dominant constant of the convergence rate is either $\mathcal{O}(\nu^{2})$ or $\mathcal{O}(c)$, which are both $n$-dependent..
@@
Under the setting of Lemma 1 and Theorem 1, for any fixed $\alpha\in[0,1]$ in the definition of $\{\alpha_{t}\}_{t\in\mathbb{N}}$ in Lemma 1, the convergence rate of Algorithm 1 is $\mathcal{O}(1/T^{\alpha+1})$, the per-iteration IFO call in expectation is $\mathcal{O}(b+n/(bt^{1-\alpha})+n/t)$ for all $t\geq 17$. Without loss of generality, assume that the target accuracy $\epsilon\in(0,1)$, define $\hat{\alpha}=\max\{\log(2)/\log(\lceil 1/\epsilon\rceil),\log(2)/\log(n)\}$, then the IFO call in expectation needed by Algorithm 1 to reach an $\epsilon$-accurate solution is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\min\left\{\log(n),\log(1/\epsilon)\right\}\right),&\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right],\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\left(\frac{n}{b}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\right)\min\left\{\log(n),\log(1/\epsilon)\right\}\right),&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right].\end{cases}$ (3)
@@
Note that the sequence $\{\mathbb{E}[F(y_{t})]\}_{t}$ also converges to $F(w^{*})$ with convergence rate $\mathcal{O}(1/t^{2\alpha})$, and when $\alpha=1$, it has the same convergence rate with the sequence $\{\mathbb{E}[F(w_{t})]\}_{t}$.
@@
In the finite-sum optimization, one of the most fundamental topics is the trade-off between the convergence rate and the per-iteration IFO cost, since essentially it is these two factors that determines the theoretical complexity of an optimization method. Handling this trade-off typically involves using different gradient estimators, using different optimization techniques, and even changing the algorithmic structure. For example, using full gradient gives the proximal gradient descent (PGD) a fast iteration convergence rate and high per-iteration IFO cost. While using the SGD estimator gives the proximal stochastic gradient descent (PSGD) contradictory characteristics. On the algorithmic structure side, note that the majority of variance reduction methods are nested, while PGD and PSGD are single-loop methods.
However, based on the proposed $p_{t}-\alpha_{t}$ correspondence, we can achieve this trade-off under the unified framework of Algorithm 1, by changing the growth rate of the $\{\alpha_{t}\}_{t\in\mathbb{N}}$ sequence. Concretely, when the acceleration strength (determined by $\alpha$) increases, the convergence rate $\mathcal{O}(1/T^{\alpha+1})$ becomes faster, and the per-iteration cost $\mathcal{O}(b+n/(bt^{1-\alpha})+n/t)$ increases at the same time.
@@
As mentioned above, less aggressive momentum acceleration (i.e., smaller $\alpha$) accompanies with smaller checkpoint update probability (i.e., smaller $p_{t}$), hence slower convergence rate can be compensated with lower per-iteration IFO cost in Algorithm 1, showing a novel continuous trade-off between the convergence rate and the per-iteration cost. The key parameter $\alpha\in[0,1]$ is thus a weight that controls the proposed acceleration-variance reduction trade-off, through which $\alpha$ essentially controls the trade-off between the convergence rate and the per-iteration cost. Two favourable consequences of this mechanism are listed as follows.
@@
First, Algorithm 1 recovers some existing IFO complexity of very different algorithms under a unified framework. Consider the $b=1$ case. When $\alpha=0$ (i.e., $\alpha_{t}\equiv 6$ for all $t\in\mathbb{N}$), the complexity of Algorithm 1 is $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$, recovering the complexity of SVRG++ while removing some elusive requirements concerning the knowledge of the solution $x^{*}$ needed by SVRG++. When $\alpha=1$ (i.e., $\alpha_{t}=\mathcal{O}(t)$), the complexity of Algorithm 1 is $\mathcal{O}(n+n/\sqrt{\epsilon})$, recovering the complexity of FISTA. Note that SVRG++ is a nested algorithm, uses the SVRG estimator, and employs the doubling technique, while FISTA is a single-loop algorithm, uses the full negative gradient as the iteration direction, and employs the Nesterov acceleration technique.
Second, we discover a series of novel complexity due to the fact that $\alpha\in[0,1]$ and can vary continuously. Consider the scenario where $n$ in (1) is very large or the computational resources are shorted, then the mini-batch size $b$ may be restricted to a small number. We show that when the target accuracy $\epsilon$ is prefixed and known, Algorithm 1 can achieve near-optimal complexity by choosing suitable $\alpha$ according to the knowledge of $n$ and $\epsilon$, under the restricted $b$ setting and regardless of the magnitude of $\epsilon$.
@@
Consider the $b=1$ case for Algorithm 1. When $n=10^{4},\epsilon=10^{-12}$, the complexity of SVRG++, FISTA, Algorithm 1 ($\alpha=1/2$), and the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ are $\mathcal{O}(10^{12})$, $\mathcal{O}(10^{10})$, $\mathcal{O}(10^{8})$ and $\mathcal{O}(10^{8})$, respectively. In fact, given $n$ and $\epsilon\in(0,1)$ satisfying
@@
Hence the complexity in expectation of Algorithm 1 matches the theoretical lower bound up to a logarithmic factor under the $\alpha$ setting (5). This observation can be extended and summarised in the following corollary.
@@
Under the setting and assumptions of Theorem 2, when the condition (4) holds, by choosing $b=1$ and $\alpha$ satisfying (5), the complexity in expectation of Algorithm 1 is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. When the condition (4) does not hold, by choosing $b=1$ and $\alpha=0$, the complexity of Algorithm 1 is $\tilde{\mathcal{O}}(n)$. Hence, Algorithm 1 achieves near-optimal complexity in both cases under the restricted $b=1$ setting.
@@
Since $\Delta_{1}<\Delta_{2}$ and $\Delta_{2}>0$ under the condition (4) and the appropriate $C_{2}$ choice described in the footnote, the interval $\Delta$ is well-defined, and the intervals in (5) are well-defined and contained in $[0,1]$. Direct calculation shows that $\alpha\geq\Delta_{1}$ implies $1/\epsilon^{1/(\alpha+1)}\leq C_{1}\sqrt{n}/\sqrt{\epsilon}$ and $\alpha\leq\Delta_{2}$ implies $n/\epsilon^{\alpha/(\alpha+1)}\leq C_{2}\sqrt{n}/\sqrt{\epsilon}$. If $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$, under the first choice in (5), we have $\mathbb{E}C_{\text{comp}}$ takes the first form in (3) and $1/\epsilon^{1/(\alpha+1)}\leq\sqrt{n}/\sqrt{\epsilon}$, hence $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$. Similarly, one sees that if $\Delta_{2}>\min\left\{\hat{\alpha},1/10\right\}$, we have $\mathbb{E}C_{\text{comp}}=\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under the second $\alpha$ choice in (5). In all, under the setting of this corollary, $\alpha$ satisfying (5) always exists and such choice of $\alpha$ yields the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
When $b=1$ and $\alpha=0$, by Theorem 2, the complexity of Algorithm 1 is $\mathcal{O}(1/\epsilon+n\log(1/\epsilon))$, which is also $\mathcal{O}(n+n\log(n))$ when $n\geq 1/\epsilon$. Under the condition $n\geq 1/\epsilon$, the first term in the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ is dominant, hence the complexity of Algorithm 1 matches the lower bound up to a logarithmic factor in this case. ∎
@@
For ease of discussion, we take $C_{2}=1$ in Corollary 1, assume that $n$ in (1) is given and focus on the $n<1/\epsilon$ case. Some comments on Corollary 1 are listed as follows. First, we give intuitive explanations for the choice (5). The largest possible interval in (5) under the $\Delta_{2}\leq\min\left\{\hat{\alpha},1/10\right\}$ case is $\left[0,\min\left\{\hat{\alpha},1/10\right\}\right]$, and the largest possible interval in (5) under the other case is $\left[\min\left\{\hat{\alpha},1/10\right\},1\right]$, which is separated from the former interval by $\min\left\{\hat{\alpha},1/10\right\}$. Since $n$ is given and $C_{2}=1$, $\Delta_{2}$ is smaller when $\log(n)/\log(1/\epsilon)$ is larger, i.e., $\epsilon$ is larger. Hence the selection rule (5) essentially suggests using mild acceleration and small checkpoint update probability when the target accuracy is low (i.e., $\epsilon$ is relatively large), and using aggressive acceleration and large checkpoint update probability when the target accuracy is high.
To understand the strategy, one may begin by investigating Theorem 2. The convergence rate of $\mathbb{E}[F(w_{T})-F(x^{*})]$ is $\mathcal{O}(1/T^{\alpha+1})$, hence it requires $\mathcal{O}(1/\epsilon^{1/(\alpha+1)})$ iterations to reach an $\epsilon$-accurate solution, which constitutes the first part of the complexity in (3) when neglecting the logarithmic factor. The second part of the complexity is $n/\epsilon^{\alpha/(\alpha+1)}$, which comes from computing full gradient at checkpoints. Note that as $\alpha$ increases, although the convergence rate improves and $1/\epsilon^{\alpha+1}$ decreases, but $n/\epsilon^{\alpha/(\alpha+1)}$ increases due to the increase of the checkpoint update probability $p_{t}$. Since $n/\epsilon^{\alpha/(\alpha+1)}$ contains a large factor $n$, it is desirable to have lower per-iteration cost in expectation (related to $p_{t}$) while maintaining the convergence rate not too slow.
The ideal trade-off between the convergence rate and the per-iteration cost depends on the relative scale between $n$ and $1/\epsilon$. In the low accuracy regime where $\epsilon$ is large, the algorithm should have relatively low per-iteration cost and can tolerate relatively slow convergence rate, and vise versa in the high accuracy regime where $\epsilon$ is small.